We are now one month into on-sky commissioning with LSSTCam.
As reported last week, there was an anomaly in the LSSTCam focal plane cooling system on 8 May. The regular maintenance schedule has been updated, based on this experience. The system was restarted and has since been functioning normally. The team is also investigating the possibility that environmental factors contributed to the anomaly, along with mitigations.
Once LSSTCam was back in normal operation, on-sky commissioning of the AOS system resumed, along with a suite of in-dome activities to commission the calibration systems. On the AOS side, the team completed a set of on-sky measurements to refine the open loop control system, including scans in elevation to verify the elevation look-up table (LUT), and then proceeded to scans of the full elevation and azimuth range of the Simonyi telescope to refine the pointing model. The team completed the set of sensitivity matrix measurements for all optical degrees of freedom. The sensitivity matrix measurements provide a translation, for each individual degree of freedom, a mapping between applied offsets and changes of the optical wavefront. In parallel, the team has been tuning the gains of the AOS closed loop control system to achieve more stable convergence in the presence of variable environmental conditions, and investigating degeneracies between optical degrees of freedom. As with during the ComCam on-sky commissioning campaign, the most challenging optical contributions to suppress have been the 1st and 2nd order spherical aberrations, which are hypothesized to be related to temperature gradients across the primary-tertiary (M1M3) and secondary (M2) mirrors. The team is analyzing telemetry data on the mirror temperatures, correlating with delivered image quality, and is designing a set of tests to evaluate and mitigate the effects of thermal gradients on the optics.
The team continued investigations of stray and scattered light this week. One test involved rastering a bright star, Alpha Centauri, across the focal plane and to slightly off-axis angles, to help model light paths through the system and to identify reflective surfaces on the telescope and dome enclosure. The full moon also presented opportunities to characterize stray light features as a function of lunar separation angle. Throughout the past month, the team has cataloged various stray and scattered light patterns that occasionally appear on some parts of the focal plane. Modeling efforts are ongoing. The light/wind screen scheduled to be installed later this year will increase baffling around the dome shutter aperture and help to minimize stray light susceptibility.
With the bright moon and some cloudy nights this week, the team made good use of the available time for in-dome calibration work. Work continues on the fine alignment of the flat field projector system. The team acquired broadband LED flats in the multiple bands, including the first measurements with the z and y filters. The broadband flats are used to correct high spatial frequency variations in the response across the focal plane. The team also took the first photon transfer curve (PTC) measurements with LSSTCam installed on the TMA that will be used to verify the readout noise, gain conversions, and saturation levels, and compare these quantities to prior electro-optical lab measurements. Additional masks specific to the LSSTCam focal plane layout were installed for the collimated beam projector to allow simultaneous measurements of the quantum efficiency and other parameters across all detectors on the focal plane.
No on-sky observations for Science Pipelines commissioning using the Feature Based Scheduler (FBS) were taken this week. Those tests are planned to resume next week. Still, Data Production has been busy generating coadds and running trial difference imaging with the compute resources at USDF using the thousands of on-sky visits acquired to date, making frequent updates to algorithms and pipeline configurations as our characterization of LSSTCam data rapidly advances, e.g., optimizing routines for instrument signature removal (ISR) and background modeling.
The currently installed filter set is grizy. Starting next week as we enter dark time, the team plans to begin the regular cycle of swapping u and y filters each lunation.